### v2

@@ Abstract
The continuous advancement of large language models (LLMs) has brought increasing attention to the critical issue of developing fair and reliable methods for evaluating their performance. Particularly, the emergence of subjective or non-subjective cheating phenomena, such as test set leakage and prompt format overfitting, poses significant challenges to the reliable evaluation of LLMs. Since evaluation frameworks often utilize Regular Expression (RegEx) for answer extraction, some models may adjust their responses to comply with specific formats that are easily extractable by RegEx. Nevertheless, the key answer extraction module based on RegEx frequently suffers from extraction errors. This paper conducts a comprehensive analysis of the entire LLM evaluation chain, demonstrating that optimizing the key answer extraction module can improve extraction accuracy, reduce LLMs’ reliance on specific answer formats, and enhance the reliability of LLM evaluation. To address these issues, we propose xFinder, a model specifically designed for key answer extraction. As part of this process, we create a specialized dataset, the Key Answer Finder (KAF) dataset, to ensure effective model training and evaluation. Through generalization testing and evaluation in real-world scenarios, the results demonstrate that the smallest xFinder model with only 500 million parameters achieves an average answer extraction accuracy of 93.42%. In contrast, RegEx accuracy in the best evaluation framework is 74.38%. xFinder exhibits stronger robustness and higher accuracy compared to existing evaluation frameworks. All resources for xFinder are available at https://github.com/IAAR-Shanghai/xFinder.
@@ 1 Introduction
Large language models (LLMs) have seen rapid advancements in recent years [1, 2]. To explore the strengths, weaknesses, and differences of various LLMs, numerous researchers have introduced a variety of evaluation benchmarks aimed at evaluating the performance of LLMs across various dimensions [3, 4, 5, 6]. Furthermore, to facilitate comparisons of different LLMs on these benchmarks, several institutions have developed unified LLM evaluation frameworks, such as LM Eval Harness [^1] LM Eval Harness is the backend for Hugging Face’s popular Open LLM Leaderboard [7]. [8] and OpenCompass [9]. These frameworks are capable of uniformly evaluating LLM in reasoning, text generation, and domain-specific knowledge.
Despite the widespread application of the aforementioned evaluation benchmarks and frameworks, some researchers have raised concerns about the fairness and reliability of current LLM evaluation efforts [10, 11]. For instance, the benchmarks may suffer from data contamination and leakage in their test sets [12, 13]. Furthermore, a systematic review has revealed multiple unreliabilities at various stages of the existing evaluation pipeline, as illustrated in Figure 1.
@@
Given the uncertainty of LLM output content, evaluation frameworks often transform datasets to achieve uniform evaluation across different tasks. For example, they may convert datasets into the Multiple-Choice Question-Answer (MCQA) with alphabet options. However, it has been shown that for some tasks, the evaluation results under this MCQA form with alphabet options may differ from those under the open-ended setting  [14].
@@
Various institutions may employ different configurations (such as temperature settings, decoding strategies, etc.) and prompts when evaluating LLMs. LLMs are highly sensitive to these settings [12]. Even minor modifications (such as changing the case of letters in options or adding extra spaces) can significantly alter the model’s responses [14, 15, 16, 17, 18], thus affecting the fairness of the evaluation outcomes. Moreover, while employing consistent prompts may ensure a degree of fairness, uniform prompts may not suit all LLMs, potentially contradicting the initial intent of evaluating the models’ inherent capabilities.
@@
Current evaluation frameworks predominantly rely on Regular Expression (RegEx) to extract answers from model outputs. For example, they may select content following phrases like "The answer is" as the key response. However, many LLMs fail to generate standardized answers [19], which complicates the extraction of key answers from their outputs. Figure 2 illustrates instances where current evaluation frameworks failed to extract responses correctly. There are two primary scenarios where regular rules fail to extract answers: first, when the model does not produce a relevant response to the evaluation question, and second, when the model’s response does not conform to standards. In the first scenario, it is unreasonable to classify a model’s failure to provide a relevant response as an error, as the evaluation aims to measure the model’s knowledge comprehension and reasoning capabilities, not merely its instruction-following ability. In the second scenario, although the LLM’s response effectively addresses the question, the response’s non-standard format impedes correct extraction using RegEx, resulting in ’false judgments’.
@@
Consequently, existing LLM evaluation practices still face numerous deficiencies in terms of fairness and reliability. How to conduct reliable evaluations has become an urgent issue in the field of LLMs. For extracting and matching key answers, the most effective and reliable methods involve using advanced LLMs (such as GPT-4) alongside manual evaluation [20]. However, due to time and financial constraints, these methods are seldom applied in practical evaluations. As a result, some initiatives have developed models, known as Judge models, through fine-tuning strategies to automatically determine the correctness of LLM responses [21, 22]. Unfortunately, these methods significantly lag behind advanced LLMs or manual evaluation in terms of generalizability.
To address the conflict between fairness and cost, we propose xFinder, a novel evaluation component designed to replace the answer extraction phase of existing frameworks. We have optimized the logic for matching extraction results, thereby enhancing the reliability of evaluations. The core of xFinder involves constructing a large-scale LLM response evaluation dataset and proposing a high-precision, low-parameter model based on this dataset to serve as the evaluator. By replacing the existing RegEx-based answer evaluation methods, xFinder can evaluate LLM outputs more accurately and efficiently, further enhancing the reliability of LLM evaluations.
To construct xFinder, we initially created the Key Answer Finder (KAF) Dataset. The training set is used to fine-tune xFinder, and we successfully fine-tuned 19 LLMs of varying sizes, all of which performed well on both the test and generalization sets. For instance, on the testing set, xFinder-qwen1505 achieved an average extraction accuracy of 96.88%. To ensure the generalizability of xFinder, we constructed a generalization set using outputs from various models on different tasks beyond the original KAF dataset. xFinder-qwen1505 achieved an average extraction accuracy of 93.42% on this dataset, demonstrating the model’s substantial generalizability. Finally, to ensure consistent experimental settings, we utilized xFinder and the traditional evaluation framework’s RegEx answer extraction module to assess 10 LLMs on a common set of evaluation tasks under real-world conditions, subsequently ranking the 10 LLMs’ performance. We found that the RegEx-based key answer extraction components in existing frameworks often fail to extract answers correctly or at all. Moreover, rankings produced by different evaluation frameworks show significant inconsistencies. In contrast, our fine-tuned xFinder models of various sizes achieved high consistency in LLM rankings, demonstrating greater stability in the evaluation results.
@@
  • We introduce xFinder, a model specifically designed for key answer extraction. The KAF dataset supports its effective training and evaluation.
  • In our extensive experiments, we demonstrate that RegEx-based evaluation methods are unreliable, while our xFinder model significantly improves reliability.
@@ 2 Related Work
From the perspective of LLM capabilities, current evaluations can be categorized into knowledge and capability evaluations [23], vertical domain evaluations [24], and alignment and safety evaluations [25, 26]. When considering the types of questions in benchmark datasets, evaluations can be divided into objective questions (such as multiple-choice, classification, mathematical questions, fill-in-the-blank, etc.) and subjective questions (such as summarization, text continuation, open-ended questions, and code generation). For objective questions, metrics like accuracy and F1-score are widely used. These metrics rely on the ability to exactly match the LLM’s output with the correct answer.
Currently, many popular unified LLM evaluation frameworks exist, including LM Eval Harness [8], OpenCompass [9], OpenAI Evals [27], and UltraEval [28]. When using objective question datasets to evaluate LLMs, these frameworks first conduct data item transformation, converting classification and fill-in-the-blank questions into multiple-choice questions. They then extract and match key answers using RegEx methods. This process either involves literally extracting the option letters for text response evaluation or extracting the probabilities of different option letters during generation for first-token evaluation. However, both methods have notable drawbacks when evaluating objective questions. In text response evaluation, inaccurately extracting the key answer from the LLM’s response can lead to incorrect evaluation results. On the other hand, first-token evaluation may not accurately reflect the LLM’s intended answer, as it overlooks the LLM’s reasoning process [29].
In addition to RegEx-based answer extraction, recent research has increasingly focused on fine-tuning LLMs for evaluation, known as LM-based methods. Specifically, these methods include training evaluation models to select the better response to the same question (Pairwise Selection) [21, 22, 30] and training evaluation models to score a given question directly, response, and reference (Pointwise Grading) [31, 30, 32]. However, these methods have significant drawbacks. The current prompt and fine-tuning dataset designs often result in evaluation LLMs with poor generalization, making them difficult to compare with strong LLMs like GPT-4 [33, 20].
Recently, the concept of reliable evaluation has been proposed  [^2] https://sites.google.com/view/real-info2024/overview. This concept aims to more scientifically assess the authenticity of content generated by LLMs and the potential risks to information systems [34]. However, systematic research and efforts to improve reliable evaluation are still lacking. To the best of our knowledge, our work is the first to conduct a systematic study in this area.
@@ 3 Problem Definition
The existing methods described in Section 2 aim to align the scores assigned by evaluation models with human-expected scores. In other words, these methods strive to ensure that the evaluation model’s judgments of LLM outputs match human judgments. Since scoring is subjective, unlike the objective determination of correctness, these methods are more suitable for open-ended tasks like text translation and summarization. In contrast, our method aims to align xFinder’s extraction of key answers from LLM outputs with the key answers expected by humans. Fundamentally, this method is more suitable for tasks with deterministic answers such as multiple-choice questions and mathematical problems. We define the key answer extraction task as follows:
Let $\mathcal{D}$ denote an evaluation dataset consisting of triplets: a question, a set of options, and the correct answer, as shown in Equation 1. Here, $q$ denotes the question, $\mathcal{C}$ denotes the set of options, and $a$ denotes the correct answer. Let $\Sigma$ be the set of all tokens, and $\mathcal{P}$ the power set.
$\mathcal{D}\subset\{(\boldsymbol{q},\mathcal{C},\boldsymbol{a})\mid\boldsymbol{q}\in\Sigma^{*},\boldsymbol{a}\in\Sigma^{*},\mathcal{C}\in\mathcal{P}(\Sigma^{*})\}$ (1)
Each data point $(\boldsymbol{q},\mathcal{C},\boldsymbol{a})$ in $\mathcal{D}$ is fed into the LLM, producing an output $\boldsymbol{y}$, as shown in Equation 2:
$\boldsymbol{y}=\text{LLM}(\boldsymbol{q},\mathcal{C},\boldsymbol{a})$ (2)
If a unique substring in $\boldsymbol{y}$ can be identified as one of the three types—Direct, Prompt wrapped, or Converted question wrapped—then this substring is considered the Key Answer Sentence $\boldsymbol{s}$. Intuitively, $\boldsymbol{s}$ is a sentence that contains the final answer to the question $\boldsymbol{q}$. The definitions of these three types and their corresponding sets are as follows:
@@
Here, $\boldsymbol{s}$ consists of relevant prompts combined with the final answer, as detailed in Equation 4. The subset $\mathcal{F}\subset\mathcal{P}(\Sigma^{*})$ represents a collection of prompts guiding to the answer, where the placeholder <final answer> indicates the position of the final answer. The expression $\boldsymbol{s}\circ\boldsymbol{f}$ denotes the new sentence formed by replacing the <final answer> token in $\boldsymbol{f}$ with $\boldsymbol{s}$.
@@
Particularly, when multiple substrings meet the criteria, it is necessary to determine if a Chain of Thought (CoT) process [^3] CoT (Chain of Thought [35]) refers to the process of progressively reasoning through $\boldsymbol{q}$ to arrive at the final result. exists within $\boldsymbol{y}$. If a CoT process is present in $\boldsymbol{y}$, $\boldsymbol{s}$ is defined as the unique substring that meets the criteria and appears after the CoT process in $\boldsymbol{y}$. If no CoT process is present, $\boldsymbol{s}$ does not exist in $\boldsymbol{y}$.
$\mathcal{S}_{1}$ =$=\{\boldsymbol{c},\tau(\boldsymbol{c})\mid\boldsymbol{c}\in\mathcal{C}\}$ (3)
$\mathcal{S}_{2}$ =$=\{\boldsymbol{s}\circ\boldsymbol{f}\mid\boldsymbol{s}\in\mathcal{S}_{1},\boldsymbol{f}\in\mathcal{F}\}$ (4)
$\mathcal{S}_{3}$ =$=\{\boldsymbol{s}\circ\boldsymbol{x}\mid\boldsymbol{s}\in\mathcal{S}_{1},\boldsymbol{x}\in\zeta(\boldsymbol{q})\}$ (5)
The definition of the key answer $\boldsymbol{k}$ is divided into two scenarios:
@@
    If $\boldsymbol{s}$ does not exist, $\boldsymbol{k}$ is set to [No valid answer];
@@
    If $\boldsymbol{s}$ exists, $\boldsymbol{k}$ is the option within the set $\mathcal{C}$ that best corresponds to what is described by $\boldsymbol{s}$, as defined in Equation 6:

### v3

@@ Abstract
The continuous advancement of large language models (LLMs) has brought increasing attention to the critical issue of developing fair and reliable methods for evaluating their performance. Particularly, the emergence of cheating phenomena, such as test set leakage and prompt format overfitting, poses significant challenges to the reliable evaluation of LLMs. As evaluation frameworks commonly use Regular Expression (RegEx) for answer extraction, models may adjust their responses to fit formats easily handled by RegEx. Nevertheless, the key answer extraction module based on RegEx frequently suffers from extraction errors. Furthermore, recent studies proposing fine-tuned LLMs as judge models for automated evaluation face challenges in terms of generalization ability and fairness. This paper comprehensively analyzes the entire LLM evaluation chain and demonstrates that optimizing the key answer extraction module improves extraction accuracy and enhances evaluation reliability. Our findings suggest that improving the key answer extraction module can lead to higher judgment accuracy and improved evaluation efficiency compared to the judge models. To address these issues, we propose xFinder, a novel evaluator for answer extraction and matching in LLM evaluation. As part of this process, we create a specialized dataset, the Key Answer Finder (KAF) dataset, to ensure effective model training and evaluation. Generalization tests and real-world evaluations show that the smallest xFinder model, with only 500 million parameters, achieves an average extraction accuracy of 93.42%. In contrast, RegEx accuracy in the best evaluation framework is 74.38%. The final judgment accuracy of xFinder reaches 97.61%, outperforming existing evaluation frameworks and judge models. All resources for xFinder are available at https://github.com/IAAR-Shanghai/xFinder.
@@ 1 Introduction
Large language models (LLMs) have seen rapid advancements in recent years (Brown et al., 2020; Achiam et al., 2023). To explore the strengths, weaknesses, and differences of various LLMs, numerous researchers have introduced a variety of evaluation benchmarks aimed at evaluating the performance of LLMs across various dimensions (Talmor et al., 2019; Clark et al., 2018; Hendrycks et al., 2020; Cobbe et al., 2021). Furthermore, to facilitate comparisons of different LLMs on these benchmarks, several institutions have developed unified LLM evaluation frameworks, such as LM Eval Harness [^1] LM Eval Harness is the backend for Hugging Face’s Open LLM Leaderboard (Beeching et al., 2023). (Gao et al., 2021) and OpenCompass (OpenMMLab, 2023). These frameworks are capable of uniformly evaluating LLM in reasoning, text generation, and domain-specific knowledge.
Despite the widespread application of the aforementioned evaluation benchmarks and frameworks, some researchers have raised concerns about the fairness and reliability of current LLM evaluation efforts (Aiyappa et al., 2023; Ye et al., 2024; Yu et al., 2024). For instance, the benchmarks may suffer from data contamination and leakage in their test sets (Zhou et al., 2023; Oren et al., 2023). Furthermore, a systematic review has revealed multiple unreliabilities at various stages of the existing evaluation pipeline, as illustrated in Figure 1.
@@
Given the uncertainty of LLM output content, evaluation frameworks often transform datasets to achieve uniform evaluation across different tasks. For example, they may convert datasets into the Multiple-Choice Question-Answer (MCQA) with alphabet options. However, it has been shown that for some tasks, the evaluation results under this MCQA form with alphabet options may differ from those under the open-ended setting  (Röttger et al., 2024).
@@
Various institutions may employ different configurations (such as temperature settings, decoding strategies, etc.) and prompts when evaluating LLMs. LLMs are highly sensitive to these settings (Zhou et al., 2023; Zheng et al., 2025). Even minor modifications (such as changing the case of letters in options or adding extra spaces) can significantly alter the model’s responses (Röttger et al., 2024; Gu et al., 2023; Sun et al., 2023; Yukun et al., 2024; Chen et al., 2024), thus affecting the fairness of the evaluation outcomes. Moreover, while employing consistent prompts may ensure a degree of fairness, uniform prompts may not suit all LLMs, potentially contradicting the initial intent of evaluating the models’ inherent capabilities.
@@
Current evaluation frameworks predominantly rely on Regular Expression (RegEx) to extract answers from model responses. For example, they often select content following phrases like “The answer is” as the key response. However, many LLMs fail to generate standardized answers (Asai et al., 2023), complicating key answer extraction from their responses. Figure 2 illustrates instances where current evaluation frameworks failed to extract responses correctly. There are two primary scenarios where regular rules fail: first, when the model’s response is irrelevant to the evaluation question, and second, when the response format does not conform to standards. In the first scenario, it is unreasonable to classify a model’s failure to provide a relevant response as an error, as the evaluation aims to measure the model’s knowledge comprehension and reasoning capabilities, not merely its instruction-following ability. In the second scenario, although the LLM’s response effectively addresses the question, the response’s non-standard format impedes correct extraction using RegEx, resulting in false judgments.
@@
In summary, existing LLM evaluation practices continue to face deficiencies in fairness and reliability, making reliable evaluations an urgent issue in the field of LLMs. The most effective methods for extracting and matching key answers use advanced LLMs (such as GPT-4) with manual evaluation (Huang et al., 2024). However, due to time and financial constraints, these methods are seldom applied in practical evaluations. As a result, some initiatives have developed fine-tuned judge models to automatically evaluate the LLM responses (Wang et al., 2023b; Zhu et al., 2023). Unfortunately, these methods significantly lag behind advanced LLMs or manual evaluations in generalizability.
To address the conflict between fairness and cost in LLM evaluation, we propose xFinder, a novel evaluator for answer extraction and matching. xFinder achieves higher judgment accuracy than RegEx-based methods and judge models, thereby improving evaluation reliability. The core of xFinder is constructing a large-scale LLM response evaluation dataset and developing a precise, efficient model to serve as the evaluator. By replacing RegEx-based answer evaluation methods, xFinder assesses LLM responses more accurately and efficiently, further enhancing evaluation reliability.
To construct xFinder, we initially created the Key Answer Finder (KAF) Dataset. We fine-tuned 19 LLMs of varying sizes using the training set, and xFinder performed well on both the test and generalization sets. For instance, xFinder-qwen1505 achieved an average extraction accuracy of 93.42% on the generalization set, demonstrating the model’s substantial generalizability. Subsequently, we compared the judgment accuracy of xFinder with mainstream evaluation frameworks and LLM-based judge models (PandaLM, JudgeLM, and GPT-4) on the generalization set. The results show that xFinder achieves a judgment accuracy of 97.61%, significantly outperforming PandaLM (51.9%), JudgeLM (78.13% for the 33B model), and GPT-4 (84.2%). Finally, we used xFinder and a traditional RegEx extraction module to assess 10 LLMs on a common set of real-world tasks and ranked their performance. We found that RegEx-based answer extraction components in existing frameworks often fail to extract answers correctly or fail entirely. Moreover, rankings from different evaluation frameworks exhibit significant inconsistencies. In contrast, our fine-tuned xFinder models consistently produced stable LLM rankings, demonstrating greater evaluation reliability.
@@
  • We introduce xFinder, a novel evaluator for answer extraction and matching in LLM evaluation, and construct the KAF dataset to support its training and testing.
  • Our experiments show that both RegEx-based frameworks and automated judge models exhibit unreliability, leading to low judgment accuracy. In contrast, our method achieves higher accuracy and improved evaluation efficiency.
@@ 2 Related Work
From the perspective of LLM capabilities, current evaluations can be categorized into knowledge and capability evaluations (Lyu et al., 2024), vertical domain evaluations (Liang et al., 2024a), and alignment and safety evaluations (Li et al., 2024; Guo et al., 2023). When considering the types of questions in benchmark datasets, evaluations can be divided into objective questions and subjective questions. For objective questions, metrics like accuracy and F1-score are widely used. These metrics rely on the ability to exactly match the LLM’s response with the correct answer.
Currently, many popular unified LLM evaluation frameworks exist, including LM Eval Harness (Gao et al., 2021), OpenCompass (OpenMMLab, 2023), UltraEval (He et al., 2024), and OpenAI Evals (OpenAI, 2024). When using objective question datasets to evaluate LLMs, these frameworks first conduct data item transformation, converting classification and fill-in-the-blank questions into multiple-choice questions. They then extract and match key answers using RegEx methods. This process involves either extracting option letters for text response evaluation or calculating the probabilities of option letters during generation for first-token evaluation. However, both methods have notable drawbacks when evaluating objective questions. In text response evaluation, inaccurately extracting the key answer from the LLM’s response can lead to incorrect evaluation results. On the other hand, first-token evaluation may not accurately reflect the LLM’s intended answer, as it overlooks the LLM’s reasoning process (Wang et al., 2024c).
In addition to RegEx-based answer extraction, recent research increasingly focuses on fine-tuning LLMs for evaluation, known as judge models. Specifically, these methods include training evaluation models to select the better response to the same question (Pairwise Selection) (Wang et al., 2023b; Zhu et al., 2023; Li et al., 2023), and training models to directly score a given question, response, and reference (Pointwise Grading) (Zheng et al., 2024; Li et al., 2023; Kim et al., 2023). However, these judge models demonstrate poor generalization when evaluating different LLMs and benchmarks (Laskar et al., 2023; Huang et al., 2024). In open-domain QA tasks, recent work has introduced the EVOUNA dataset, which can be used to evaluate the accuracy of automated evaluation methods (Wang et al., 2024a).
Recently, the concept of reliable evaluation has been proposed [^2] https://sites.google.com/view/real-info2024/overview. This concept aims to rigorously assess the authenticity of content generated by LLMs and the potential risks to information systems (Augenstein et al., 2023). However, systematic research and efforts to improve reliable evaluation are still lacking. To the best of our knowledge, our work is the first to conduct a systematic study in this area.
@@ 3 Problem Definition
The existing methods described in Section 2 aim to align the scores assigned by evaluation models with human-expected scores. In other words, these methods strive to ensure that the evaluation model’s judgments of LLM outputs match human judgments. Since scoring is subjective, unlike the objective determination of correctness, these methods are more suitable for open-ended tasks like text translation and summarization. In contrast, our method aims to align xFinder’s extraction of key answers from LLM outputs with the key answers expected by humans. Fundamentally, this method is more suitable for tasks with deterministic answers, such as multiple-choice questions and mathematical problems. We define the key answer extraction task as follows:
Let $\mathcal{D}$ denote an evaluation dataset consisting of triplets: a question, a set of options, and the correct answer, as shown in Equation 1. Here, $q$ denotes the question, $\mathcal{C}$ denotes the set of options, and $a$ denotes the correct answer. Let $\Sigma$ be the set of all tokens and $\mathcal{P}$ the power set.
$\mathcal{D}\subset\{(\bm{q},\mathcal{C},\bm{a})\mid\bm{q}\in\Sigma^{*},\bm{a}\in\Sigma^{*},\mathcal{C}\in\mathcal{P}(\Sigma^{*})\}$ (1)
Each data point $(\bm{q},\mathcal{C},\bm{a})$ in $\mathcal{D}$ is fed into the LLM, producing an output $\bm{y}$, as shown in Equation 2:
$\bm{y}=\text{LLM}(\bm{q},\mathcal{C},\bm{a})$ (2)
If a unique substring in $\bm{y}$ can be identified as one of the three types—Direct, Prompt wrapped, or Converted question wrapped—then this substring is considered the Key Answer Sentence $\bm{s}$. Intuitively, $\bm{s}$ is a sentence that contains the final answer to the question $\bm{q}$. The definitions of these three types and their corresponding sets are as follows:
@@
Here, $\bm{s}$ consists of relevant prompts combined with the final answer, as detailed in Equation 4. The subset $\mathcal{F}\subset\mathcal{P}(\Sigma^{*})$ represents a collection of prompts guiding to the answer, where the placeholder <final answer> indicates the position of the final answer. The expression $\bm{s}\circ\bm{f}$ denotes the new sentence formed by replacing the <final answer> token in $\bm{f}$ with $\bm{s}$.
@@
Particularly, when multiple substrings meet the criteria, it is necessary to determine if a Chain of Thought (CoT) process [^3] CoT (Chain of Thought (Wei et al., 2022)) refers to the process of progressively reasoning through $\bm{q}$ to arrive at the final result. exists within $\bm{y}$. If a CoT process is present, $\bm{s}$ is defined as the unique substring that meets the criteria and appears after the CoT process in $\bm{y}$. If no CoT process is present, $\bm{s}$ does not exist in $\bm{y}$.
$\mathcal{S}_{1}$ =$=\{\bm{c},\tau(\bm{c})\mid\bm{c}\in\mathcal{C}\}$ (3)
$\mathcal{S}_{2}$ =$=\{\bm{s}\circ\bm{f}\mid\bm{s}\in\mathcal{S}_{1},\bm{f}\in\mathcal{F}\}$ (4)
$\mathcal{S}_{3}$ =$=\{\bm{s}\circ\bm{x}\mid\bm{s}\in\mathcal{S}_{1},\bm{x}\in\zeta(\bm{q})\}$ (5)
The definition of the key answer $\bm{k}$ is divided into two scenarios:
@@
    If $\bm{s}$ does not exist, $\bm{k}$ is set to [No valid answer];
@@
    If $\bm{s}$ exists, $\bm{k}$ is the option within the set $\mathcal{C}$ that best corresponds to what is described by $\bm{s}$, as defined in Equation 6: